OpenStudy (anonymous):

Let f(x) = x – 3 and g(x) = x + 11. Find f(x) • g(x). x2 – 8x – 33 x2 – 8x + 14 x2 + 8x + 14 x2 + 8x – 33

myininaya (myininaya):

\[f(x) \cdot g(x)=(x-3)(x+11)\] Just expand and simplify

OpenStudy (anonymous):

so it would be x2-8x+14 where does the 8x come from?

myininaya (myininaya):

Nope. -3(11) is not 14

myininaya (myininaya):

Also 11x-3x is not -8x

myininaya (myininaya):

\[(x-3)(x+11)=x(x+11)-3(x+11)=x \cdot x + x \cdot 11 -3 \cdot x -3 \cdot 11\]

OpenStudy (anonymous):

f(x).g(x)= (x-3)(x+11) =x(x+11)-3(x+11) =x2+11x-3x-33 =x2+8x-33

OpenStudy (anonymous):

awesome thanks i understand